OpenStudy (anonymous):

what would the graph of the equation |x-y|=3 look like? Choices: \\ , // , / , \/

zepdrix (zepdrix):

\[\large\rm |x-y|=3\]\[\large\rm x-y=\pm3\]Solve for y,\[\large\rm y=x\mp3\]So we have two equations:\[\large\rm y=x+3, \qquad y=x-3\]Graph those to see which lines you should be getting :)

OpenStudy (anonymous):

thank you so much!